Two 1.60kg carts on a low-friction track are rigged with magnets so that the carts repel each other when they approach each other. One cart has an initial velocity of +0.323 m/s, and the other has an initial velocity of -0.147 m/s. What is the maximum energy stored in the magnetic field during the collision, assuming that the carts never touch each other?


1
Expert's answer
2021-04-22T11:06:47-0400

From the conservation of energy:



"E=K_1+K_2=0.5mv_1^2+0.5mv_2^2=0.5m(v_1^2+v_2^2)""E=0.5(1.6)(0.323^2+(-0.147)^2)=0.1007\ J."